OpenStudy (anonymous):

If R is inversely proportional to A, and R=4 when A=100, what is the value of R when A=250?

hartnn (hartnn):

inverse proportion means product is constant. so, R1A1=R2A2 4*100 = R2*250 find R2

OpenStudy (anonymous):

im confused with the R1 and A2, what are those?

hartnn (hartnn):

lets do it in another way inverse proportion means product is constant. so, AR = k (a constant) find k , given R=4, A=100

OpenStudy (anonymous):

k=400

hartnn (hartnn):

right, now u have AR = 400 what is R , when A = 250

OpenStudy (anonymous):

umm 1.6?

hartnn (hartnn):

thats your R correct :)

OpenStudy (anonymous):

thanks a lot :D

hartnn (hartnn):

welcomes a lot ^_^